OpenStudy (anonymous):

First to get correct answer gets a fan and medal! A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the rate in feet per second

hero (hero):

Hint: r = rate in feet per second = 96 and h = time to hit the ground = 0 so: 0 = 96t - 16t^2 Solve for t

OpenStudy (anonymous):

how do you solve for the t squared?

hero (hero):

0 = -16t^2 + 96t factor out -16t to get: 0 = -16t(t - 6) Finish solving for t

OpenStudy (anonymous):

how would you do that?

hero (hero):

Use the multiplication rule ab = 0 : Either a = 0, or b = 0 So either -16t = 0 or t - 6 = 0

OpenStudy (anonymous):

so then it would be 0=-16t(0) but when you multiply anything times zero your answer is zero isn't it?

hero (hero):

How do you solve t - 6 = 0 ?

OpenStudy (anonymous):

t=6

OpenStudy (anonymous):

so is that right?

hero (hero):

Yes, correct The time the ball hits the ground is 6 seconds.